OpenStudy (anonymous):

Showing all of your work, solve for x: -9x + 6 = -x + 4

OpenStudy (alexalas):

do you know how to move numbers from one side of the equation to the other?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

do you know where to start? @Mlink1290

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Ok you want to combin like terms... Do you know what that is? @Mlink1290

OpenStudy (alexalas):

alright take 5x=3x+2 to move a number from one side you subtract if positive and add if negative so 5x-3x=3x-3x+2 2x/2=2/2 x=1

OpenStudy (alexalas):

so -9x + 6 = -x + 4 move -9x to the other side please

OpenStudy (anonymous):

Step 1: Add x to both sides. −9x+6+x=−x+4+x −8x+6=4 Step 2: Subtract 6 from both sides. −8x+6−6=4−6 −8x=−2 Step 3: Divide both sides by -8. −8x/-8 =-2/-8 x=1/4 Answer: x=1/4

OpenStudy (alexalas):

@BlueishPastel don't just give the answer please do it step by step.

OpenStudy (anonymous):

I did do it step by step c;

OpenStudy (alexalas):

well Mlink1290 do you under stand?